OpenStudy (anonymous):

simplified a variable expression as much as possible (4+3)x+2x-5 ( I Got 9x-5)

OpenStudy (anonymous):

(4+3)x+2x-5 7x+2x-5 7x+2x=9x 9x-5 Is This Correct ??

OpenStudy (dobby1):

first you have add what is in the parenthases giving you 7x+2x-5

OpenStudy (anonymous):

4+3=7 so 7x+2x=9x 9x-5 So yeah you got it right:)

OpenStudy (dobby1):

yep you got it right